A stock index consists of these 4 companies: Stock D, 11 million shares, $19 per share Stock E, 103 million shares, $25 per share Stock F, 35 million shares, $44 per share Stock G, 73 million shares, $101 per share If the divisor is 5.2 million, then what is the index level now? Round your answer to the nearest point
Answer #1

1) A STOCK CONSISTS OF THESE 4 COMPANIES STOCK D 11 MILLION SHARES $ 19 PER SHARE STOCK E 103 MILLION SHARES $25 PER SHARE STOCK F 35 MILLION SHARES,$44 PER SHARE STOCK G 73 MILLION SHARES,$101 PER SHARE IF THE DIVISOR IS 5,2 MILLION,THEN WHAT IS THE INDEX LEVEL NOW? ROUND YOUR ANSWER TO NEAREST POINT?\

A) First let us check all the stocks and share values

1.stock D = 11*19=209

2.stock E =103*25=2575

3.stock F =35*44=1540

4.stock G =73*101=7373

DIVISOR IS 5.2 MILLIONS SO

209+2575+1540+7373=11697

11697/5,2=2249.42 is equivalent to 2250